OpenStudy (anonymous):

Eliminate the parameter. x = t2 + 1, y = t2 - 1

OpenStudy (primeralph):

t2 = x-1 y = x-1-1 = x-2

OpenStudy (anonymous):

what?

OpenStudy (primeralph):

Read.

OpenStudy (anonymous):

Well its kinda hard if I don't understand what I'm looking at.

OpenStudy (anonymous):

\[x=t^2+1, y=t^2-1\]

OpenStudy (primeralph):

y = x-2.

OpenStudy (anonymous):

okay...

OpenStudy (primeralph):

More like Thank You?

OpenStudy (anonymous):

No? You didn't even explain anything, that doesn't help at all. And your rude. Goodbye.

OpenStudy (primeralph):

So are you.